Directions: Read and analyze the given problems below. Write your solutions and final answer on a separate sheet of paper. 1. ILAW Manufacturing company produces bulbs that last a mean of 900 hours with a standard deviation of 110 What is the probability that the mean lifetime of a random sample of 15 of these bulbs is less than 850 hours?



SOLUTION:

1
Expert's answer
2022-04-11T17:17:51-0400

"P(\bar X<850)=P(Z<\frac{850-900}{\frac{110}{\sqrt{15}}})=P(Z<-1.76)=0.0392."